OpenStudy (anonymous):

What is the 35th term of the arithmetic sequence where a1 = 13 and a17 = –67?

OpenStudy (anonymous):

4. d=-5, a35 =?,a1=13 a35 = 13+(34)-5 a35 = 13 -170 a35 = -157,answer